# Comparison of the Effects of Interaction with Intentional Agent and Artificial Intelligence using fNIRS

### v1

@@ A Wavelet-Based Approach for decompose signal
Activity Measure: Examining Average Differences in Neural Responses between Game and Rest Modes
Our analysis began by setting up a mathematical framework to interpret and extract meaningful information from brain activity. We concentrated on the average intensity of signals, consolidating the data collected during 10-minute gameplay sessions at designated locations and frequencies. Following this, we calculated the mean value within this context. A similar procedure was applied to the 5-minute rest periods. Essentially, we combined the data from both rest and gameplay phases, computed individual averages for each, and subsequently assessed the disparity between these averages. This discrepancy served as a measurable indicator of activity, expressed as:
$\boldsymbol{Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)}$ (1)
A Wilcoxon Analysis on HbO Activation with Subsequent Paired t-test Scrutiny
In exploring the disparities in mean HbO activation between human and AI interaction conditions during an fNIRS cognitive test, utilizing the Wilcoxon signed-rank non-parametric test [67, 68] was crucial. This approach, structured as a within-subjects study, involved comparing matched pairs of HbO activation levels across both intentional agent and AI agent conditions for each participant, with the choice of non-parametric tests based on their independence from assumptions of normality in data distribution.
@@ A Wilcoxon Analysis on HbO Activation with Subsequent Paired t-test Scrutiny
Acknowledging the diversity in brain structures and responses to stimuli among individuals, and considering the relative nature of fNIRS imaging data, the Wilcoxon signed-rank test proved valuable for intra-individual comparisons. It calculated signed differences, ranked them, and examined whether these ranks deviated significantly from null hypothesis expectations [69], positing no HbO activation difference between intentional and AI agent conditions.
Shifting to the latter part, after utilizing the Wilcoxon test to identify the focal point, we scrutinized the normality of the data at that specific juncture using the Shapiro-Whisk test [70, 71]. This step was pivotal in determining the appropriateness of deploying the paired t-test [72] for a more detailed analysis. If the data met the assumptions of normality, employing the paired t-test would enable a deeper exploration of the observed distinctions in HbO activation between intentional and AI agent conditions.
@@
Our study posited the hypothesis that distinct channels within the prefrontal cortex play specialized roles in our task, showcasing diverse activity patterns. The two-dimensional points in the prefrontal cortex indeed revealed varied activation patterns, robustly affirming our initial hypothesis.
The Wilcoxon signed-rank analysis, detailed in the table below, highlights a significant effect observed at a specific point. While numerous points in the prefrontal cortex did not exhibit significant differences in activation between human and AI interactions, Channel 1 in the frequency range of 0.00978 to 0.0196 Hz emerged as particularly noteworthy(p-value=0.00262). The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels are shown in Table.1.
@@
Table 1: The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
We explored the feasibility of employing paired t-tests for the data corresponding to the identified point with a significant difference revealed by the Wilcoxon test. To assess the normality of the paired data, we conducted the Shapiro-Wilk test. Our findings indicate that we did not find sufficient evidence to reject the null hypothesis, suggesting that the data lack a clear tendency toward a specific distribution. The result of Shapiro-Wilk test that shows p-value = 0.340 for data in situation 1 and p-value = 0.542 for data in situation 2. Also, the Q-Q plot is shown in Fig.7.
@@
Figure 7: Q-Q plot for checking normality. It is aligned with Shapiro-Wilk test
Consequently, the paired t-test is deemed appropriate for conducting further in-depth analysis. Although the p-value in the one-tailed t-test increased, it still remained significantly lower than the determined significance level in comparison with the Wilcoxon test.You can find details related to this statistical test in Table.2. Additionally, for clarification of the differences in means, we utilized box plots with error bars, as shown in Fig.8. To validate these findings, we employed Bonferroni’s analysis [73] test, adjusting our significance level to account for the multiple comparisons. Given the 24 measurement points, we reduced the significance level to 0.002 by dividing 0.05 by the number of points. However, these adjusted results failed to maintain statistical significance.
Furthermore, the lack of significant differences in other channels (in this analysis) suggests that we successfully controlled the desired conditions. These areas are typically associated with inhibition [45], emotional regulation [74], and mental workload [75], as supported by previous studies.
Overall, our results shed light on the intricate neural dynamics of social interactions with intentional agents, offering valuable insights into the specialized roles of different prefrontal cortex channels and their adaptation to the complexities of human-AI interactions.
@@
Table 2: Paired t-test Results
Figure 8: Comparison of activation levels between human and AI interactions using the box plot of means considering standard errors.
@@ Discussion
Our experiment offers valuable insights into the transition from common-sense psychology to cognitive psychology, particularly regarding interactions with intentional versus non-intentional agents. By utilizing fNIRS, we demonstrate the ability to capture and analyze these nuances, thereby advancing our understanding of human interactions in real-world settings and the underlying mental states involved.
In this cognitive experiment, our primary objective is to explore the complexities of mental states elicited by interactions with intentional agents, using fNIRS to probe neural activity in ecologically valid environments. Our study unveils a distinctive neural signature within Channel 1, located between FpZ and AFZ according to the 10-20 system, across the frequency range of 0.00978 to 0.0196 Hz.
Our findings are consistent with prior observations using more spatially precise neuroimaging techniques like fMRI and PET [32, 31, 36], affirming the robustness of our results across methodologies. The BioImage Suite website  [^2] https://bioimagesuiteweb.github.io/webapp/ is used to generate an approximate graphical representation of the brain region illustrating in Fig.9.
@@
Leveraging the temporal resolution advantages of fNIRS and wavelet analysis, we go deeper into the underlying brain mechanisms responsible for these observed differences. Our results suggest a link between this neural function and a low-frequency blood flow pattern, particularly in the mental representation related to the presence of another person. Notably, we found no significant difference in this fNIRS channel at higher frequencies when applying within-subject analysis, indicating that the sense of presence manifested in longer waves is associated with more stable processes. It appears that the representation of another person’s presence is continuous, unlike tasks involving attention or decision-making, where mental states fluctuate constantly.
Furthermore, our straightforward approach to identifying channels related to cognitive load during the tennis game task involved imagining a binary space where the result of $Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)$ had to be positive, indicating a greater relative concentration during gameplay. We labeled positive results as success and negative results as failure within this binary space. The success-to-failure ratio for 24 data points was calculated, revealing a substantial success rate of 83% (in binomial test p < .00000001) within the frequency band of 0.0195-0.0392 Hz in Channel 1. This suggests that while Channel 1 is involved, different frequencies of blood flow changes are implicated in game workload and the subjective sense of interaction with intentional agents, indicating a multifunctional role for certain brain areas.
Overall, our findings regarding the difference in frequency bands contribute to bridging the gap between common-sense psychology and scientific psychology. In simpler terms, it’s akin to a valve opening blood flow with a certain concentration and minor changes related to the feeling of presence, while gameplay opens another valve that transfers a concentration with more pronounced changes to the frontal cortex. This insight enhances our understanding of the intricate connection between the mind and the brain.
However our results highlight a distinct difference between interactions with intentional agents and artificial agent, it’s imperative to acknowledge several limitations that may influence the interpretation of our findings as follows:
@@
    Our study’s sample population was limited to students aged between 18 and 30, potentially restricting the generalizability of our findings to other demographic groups.
@@
    The spatial inaccuracy of fNIRS caps based on the 10-20 channel system may have impacted the precision of our results due to variations in skull shapes among individuals.
@@
    Another important limitation of this research project was the limited number of implemented channels for measurements, while all areas of the head have the potential to receive signals. We focused solely on specific regions within the frontal cortex for reasons mentioned in previous sections.

### v2

@@ A Wavelet-Based Approach to Decompose Signal
Quantifying Brain Activity: GLM Analysis of fNIRS Data
The analysis of fNIRS data was conducted using the GLM approach [73, 74], a widely utilized statistical technique in neuroimaging studies [75, 76, 77]. Our experimental paradigm involved alternating periods of task engagement and rest, each lasting 1 minute and 30 seconds, respectively, repeated ten times over a 15-minute session. Central to this method was the construction of a design matrix tailored to our experimental paradigm. In the design matrix, we assigned a value of 0 for rest periods and 1 for task engagement periods, effectively capturing the temporal structure of the experiment (depicted in Fig.7). Eventually, mean beta values were calculated as a measure of activity using this method. Beta values represent the estimated coefficients associated with each regressor in the GLM. By averaging these beta values across participants or conditions, we obtained a measure of the average neural activity during task engagement relative to rest periods. This approach allowed us to quantify and compare the magnitude of neural responses across different experimental conditions or groups.
Figure 7: Sample image: Comparison of HbO2 and HHb activity in level 9 (0.00978-0.0196 Hz) of Daubechies wavelet. The GLM method was applied for all four preferred levels across all channels in a manner that we imposed expected activity (1) during gameplay moments and inactivity (0) during REST moments in the design matrix.
A Wilcoxon Analysis on Activation with Subsequent Paired t-test Scrutiny
In exploring the disparities in mean activation between human and AI interaction conditions during an fNIRS cognitive test, utilizing the Wilcoxon signed-rank non-parametric test was crucial. This approach, structured as a within-subjects study, involved comparing matched pairs of both oxyHb and deoxyHb activation levels across both intentional agent and AI agent conditions for each participant [78, 79]. The choice of non-parametric tests was based on their independence from assumptions of normality in data distribution.
@@ A Wilcoxon Analysis on Activation with Subsequent Paired t-test Scrutiny
Acknowledging the diversity in brain structures and responses to stimuli among individuals, and considering the relative nature of fNIRS imaging data, the Wilcoxon signed-rank test proved valuable for intra-individual comparisons. It calculated signed differences, ranked them, and examined whether these ranks deviated significantly from null hypothesis expectations [80], positing no activation difference between intentional and AI agent conditions.
Shifting to the latter part, after utilizing the Wilcoxon test to identify the focal point, we scrutinized the normality of the data at that specific juncture using the Shapiro-Wilk test [81, 82]. This step was pivotal in determining the appropriateness of deploying the paired t-test for a more detailed analysis. If the data met the assumptions of normality, employing the paired t-test [83] would enable a deeper exploration of the observed distinctions in activation between intentional and AI agent conditions.
@@
This study posited the hypothesis that distinct channels within the prefrontal cortex play specialized roles in the considered task, showcasing diverse activity patterns. The two-dimensional points in the prefrontal cortex indeed revealed varied activation patterns, robustly affirming the initial hypothesis.
The Wilcoxon signed-rank analysis, detailed in the table below, highlights a significant effect observed at a specific point. While numerous points in the prefrontal cortex did not exhibit significant differences in activation between human and AI interactions, Channel 1 in the frequency range of 0.00978 to 0.0196 Hz emerged as particularly noteworthy(oxyHb’s p-value is 0.00262 and deoxyHb’s p-value is 00.00521). The results of Wilcoxon signed-rank test for different frequency ranges and fNIRS channels are shown in Table.1 and Table.2.
@@
Table 1: The results of Wilcoxon signed-rank test of oxyHb for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
Ch1 0.786095126 0.618795321 0.711075639 0.00521158 *
Ch2 0.722298542 0.636987456 0.502983659 0.611058963
Ch3 0.540963259 0.568812375 0.772345789 0.765540802
Ch4 0.498527441 0.359874563 0.479354592 0.501219566
Ch5 0.653717553 0.192655897 0.630291578 0.405288927
Ch6 0.880574623 0.410258846 0.411058952 0.774556922
Table 2: The results of Wilcoxon signed-rank test of deoxyHb for different frequency ranges and fNIRS channels. Each column corresponds to a frequency range and each row corresponds to a fNIRS channel
The feasibility of employing paired t-tests for the data corresponding to the identified point with a significant difference revealed by the Wilcoxon test was explored. To assess the normality of the paired data, the Shapiro-Wilk test was conducted. The findings indicate that sufficient evidence was not found to reject the null hypothesis, suggesting that the data lack a clear tendency toward a specific distribution. The Shapiro-Wilk test results indicate that the p-value for oxyHb data in situation 1 is 0.340, and for situation 2 is 0.542. Similarly, for deoxyHb, the p-value is 0.411 in situation 1 and 0.512 in situation 2. These p-values suggest that the data in both situations exhibit normal distribution characteristics, as none of the values fall below the conventional significance threshold of 0.05. Also, the Q-Q plot is shown in Fig.8.
@@
Figure 8: Q-Q plot for oxyHb. It is aligned with Shapiro-Wilk test
The paired t-test is deemed appropriate for conducting further in-depth analysis. Although the p-value in the one-tailed t-test increased, it still remained significantly lower than the determined significance level in comparison with the Wilcoxon test (oxyHb’s p-value is 0.0047 and deoxyHb’s p-value is 0.0237).You can find details related to this statistical test in Table.3 for oxyHb and Table.4 for deoxyHb. For clarification of the differences in means, box plots for oxyHb with error bars was utilized, as shown in Fig.9. To validate these findings, Bonferroni’s analysis [84] test was employed, adjusting the significance level to account for the multiple comparisons. Given the 24 measurement points, the significance level was reduced to 0.002 by dividing 0.05 by the number of points. However, these adjusted results failed to maintain statistical significance.
The lack of significant differences in other channels (in this analysis) suggests that the desired conditions were successfully controlled. These areas are typically associated with inhibition [46], emotional regulation [85], and mental workload [86], as supported by previous studies.
Overall, the results shed light on the intricate neural dynamics of social interactions with intentional agents, offering valuable insights into the specialized roles of different prefrontal cortex channels and their adaptation to the complexities of human interactions.
@@
Table 3: Paired t-test results for oxyHb
95% Confidence Interval of the Difference
Mean Std. Deviation Std. Error Mean Lower Upper t df p-value
-0.03342 0.09567 0.01426 -0.06216 -0.00467 -2.343 44 0.0237
Table 4: Paired t-test results for deoxyHb
Figure 9: Comparison of activation levels between human and AI interactions using the box plot of means for oxyHb considering standard errors.
The difference in the p-value between between oxyHb and deoxyHb can be explained considering the nature of the game [87, 88, 60]. Although both statistically indicate a difference between the two states, it seems that in this experimental setting, oxyHb demonstrates this difference more clearly.
@@ Discussion
The experiment offers valuable insights into the transition from common-sense psychology to cognitive psychology, particularly regarding interactions with intentional versus non-intentional agents. By utilizing fNIRS, the ability to capture and analyze these nuances is demonstrated, thereby advancing the understanding of human interactions in real-world settings and the underlying mental states involved.
In this cognitive experiment, the primary objective is to explore the complexities of mental states elicited by interactions with intentional agents, using fNIRS to probe neural activity in ecologically valid environments. This study unveils a distinctive neural signature within Channel 1, located between FpZ and AFZ according to the 10-20 system, across the frequency range of 0.00978 to 0.0196 Hz. In other word, the content of thought have correlation with brain signals that captured through fNIRS imaging.
The findings are consistent with prior observations using more spatially precise neuroimaging techniques like fMRI and PET [33, 32, 37], affirming the robustness of the results across methodologies. The BioImage Suite website  [^2] https://bioimagesuiteweb.github.io/webapp/ is used to generate an approximate graphical representation of the brain region illustrating in Fig.10.
@@
Leveraging the temporal resolution advantages of fNIRS and wavelet analysis, deeper exploration into the underlying brain mechanisms responsible for these observed differences is undertaken. The results suggest a link between this neural function and a low-frequency blood flow pattern, particularly in the mental representation related to the presence of another person. Notably, no significant difference in this fNIRS channel at higher frequencies was found when applying within-subject analysis, indicating that the sense of presence manifested in longer waves is associated with more stable processes. It appears that the representation of another person’s presence is continuous, unlike tasks involving attention or decision-making, where mental states fluctuate constantly.
Furthermore, the employed straightforward approach to identifying channels related to cognitive load during the tennis game task involved imagining a binary space where the result of $Activity=(Mean\;of\;Total\;Gameplay\;Period)-(Mean\;of\;Total\;Rest\;Period)$ had to be positive, indicating a greater relative concentration during gameplay. Positive results were labeled as success and negative results were labeled as failure within this binary space. The success-to-failure ratio for 24 data points was calculated, revealing a substantial success rate of 83% (in binomial test p < 0.00000001) within the frequency band of 0.0195-0.0392 Hz in Channel 1. This suggests that while Channel 1 is involved, different frequencies of blood flow changes are implicated in game workload and the subjective sense of interaction with intentional agents, indicating a multifunctional role for certain brain areas.
Overall, the findings concerning differences in frequency bands serve to bridge the gap between common-sense psychology and scientific psychology. To simplify, consider it akin to regulating blood flow through a valve: one valve manages a consistent flow with subtle adjustments, impacting the sense of presence, while another valve oversees a flow with more pronounced changes, particularly affecting the frontal cortex during gameplay. Drawing on this analogy, we hypothesize that the varying speeds of different blood streams in the brain correspond to distinct mental states. This concept is elucidated by fNIRS, which provides both spatial and temporal resolution, further refined by computational techniques such as wavelet analysis. Such an approach can offer profound insights into the areas involved in various mental operations or hubs in the cerebral cortex, such as Brodmann’s Area 10 [89, 90, 91]. Nonetheless, this remains a conjecture to be explored in future research endeavors.
The results highlight a distinct difference between interactions with intentional agents and artificial agent. It’s imperative to acknowledge several limitations that may influence the interpretation of this research’s findings as follows:
@@
    This study’s sample population was limited to students aged between 18 and 30, potentially restricting the generalizability of the findings to other demographic groups.
@@
    The spatial inaccuracy of fNIRS caps based on the 10-20 channel system may have impacted the precision of the results due to variations in skull shapes among individuals.
@@
    Another important limitation of this research project was the limited number of implemented channels for measurements, while all areas of the head have the potential to receive signals. Sole focus was placed on specific regions within the frontal cortex for reasons mentioned in previous sections.